OpenStudy (anonymous):

Evaluate the integral of ln(x) dx from 1 to 2

OpenStudy (anonymous):

\[\int\limits_{1}^{2}\ln x dx=(xlnx-x)|_{1}^{2}=2\ln2-3\]

OpenStudy (anonymous):

I got 2ln2 -1

OpenStudy (nowhereman):

The latter is correct \[2\ln 2 - 2 - (1\ln 1 - 1) = 2\ln 2 - 2 + 1 = 2\ln 2 -1\]

OpenStudy (anonymous):

Ok thanks!

OpenStudy (anonymous):

intg(0 to2)lnxdx=xlnx -x]=x(lnx-x)]=2\(ln2-2)\-(1ln1-1) =2ln2 -2+1=2ln2-1